OpenStudy (ria23):

Which point lies on the graph of the function shown below? y = -x^2 + 5x - 3 How would I solve this?

OpenStudy (ria23):

I tried getting it into intercept form but it won't work, same with vertex. >.<

jimthompson5910 (jim_thompson5910):

I'm guessing you were given a list of points?

OpenStudy (ria23):

M'hmm. Did yhu want me to post 'em?

jimthompson5910 (jim_thompson5910):

yes please

jimthompson5910 (jim_thompson5910):

always post the full question, instructions and answer choices you don't always have to post the answer choices unless they are critical like in this case

jimthompson5910 (jim_thompson5910):

let's test (0,3)

jimthompson5910 (jim_thompson5910):

y = -x^2 + 5x - 3 3 = -(0)^2 + 5(0) - 3 ... plug in (x,y) = (0,3) 3 = -3 ... this is FALSE so because we get a false equation, this means (0,3) does NOT lie on y = -x^2 + 5x - 3

jimthompson5910 (jim_thompson5910):

I'll let you check the others

OpenStudy (ria23):

(2,3) ? It's the only one I plugged in and got 3 = 3